When ¼ is written with denominator as 12, its numerator is (A) 3 (B) 8 (C) 24 (D) 12

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(1 × 3)/(4 × 3) = 3/12
Consider, 3/12
Divide both numerator and denominator by 3
= 1/4
Hence, the answer is 3


Thanks